# Scaling Laws with Vocabulary: Larger Models Deserve Larger Vocabularies

### v1

@@ Abstract
Figure 1: The relationship between non-vocabulary parameters $N_{\rm nv}$ and their optimal vocabulary parameters $N_{\rm v}^{\rm opt}$ follows a power law. The scaling rate of vocabulary parameters should be slower than that of non-vocabulary parameters, expressed as $\gamma<1$. Empirical results align with predictions from all proposed approaches, with larger circles indicating higher loss values.
@@ 1 Introduction
Figure 2: Most current LLMs have suboptimal vocabulary parameters due to vocabulary sizes, which are smaller than the predicted optimal values. Among the current models, StarCoder2-3B, OLMo-7B, InternLM2-20B, and Gemma2-27B have vocabulary sizes that come closest to the optimal allocation for their respective model sizes.
@@ 4.3 Approach 3: Parametric fit of loss formula
Figure 6: The vocabulary parameters of current public LLMs and the predicted optimal vocabulary parameters by our Approach 3 (§4.3). Here we consider the practical scenarios where parameters and training data are not necessarily equally scaled. As illustrated, the vocabulary parameters remain predominantly underestimated. With the exception of Gemma2-9B, all models allocate a smaller parameter count than our predicted optimal vocabulary parameter.
@@ 6 Related work
The vocabulary of a language model influences its performance significantly [59, 68, 71]. A larger vocabulary size helps cover more words thus reducing the likelihood of out-of-vocabulary (OOV) cases [19]. Takahashi and Tanaka-Ishii [59] find that larger vocabularies are better at capturing the true statistical distribution of language. Similarly, expanding vocabulary in multilingual models [68] improves performance, especially for low-resource languages. However, large vocabularies [29] increase the computational overhead during both training and generation phases. Liao et al. [32] show that low-frequency words have few instances in the training data, leading to insufficient information to train robust representations if the vocabulary is too large. To this end, our work fills an under-explored gap: How to optimally allocate the vocabulary size?

### v2

@@ Abstract
Figure 1: The relationship between non-vocabulary parameters $N_{\rm nv}$ and the corresponding optimal vocabulary parameters $N_{\rm v}^{\rm opt}$ follows a power law, where $N_{\rm v}^{\rm opt}$ should be scaled slower than $N_{\rm nv}$ as $\gamma<1$. Empirical results align with predictions of our proposed approaches, with larger circles indicating higher loss values. Here $V$ refers to the vocabulary size i.e. the number of distinct tokens.
@@ 1 Introduction
Figure 2: Vocabulary parameters of popular LLMs and predicted optimal vocabulary parameters at a compute-optimal number of training tokens. Most current LLMs have suboptimal vocabulary parameters due to vocabulary sizes, which are smaller than the predicted optimal values. Among the current models, StarCoder2-3B, OLMo-7B, InternLM2-20B, and Gemma2-27B have vocabulary sizes that come closest to the optimal allocation for their respective model sizes.
@@ 4.3 Approach 3: Parametric fit of loss formula
Figure 6: Vocabulary parameters of popular LLMs and predicted optimal vocabulary parameters at their reported number of training tokens, as determined by our Approach 3 (§4.3). Here we consider the practical scenarios where parameters and training data are not necessarily equally scaled. As illustrated, the vocabulary parameters remain predominantly underestimated. With the exception of Gemma2-9B, all models allocate a smaller parameter count than our predicted optimal vocabulary parameter.
@@ 6 Related work
The vocabulary of a language model influences its performance significantly [62, 71, 74]. A larger vocabulary size helps cover more words thus reducing the likelihood of out-of-vocabulary (OOV) cases [21]. Takahashi and Tanaka-Ishii [62] find that larger vocabularies are better at capturing the true statistical distribution of language. Similarly, expanding vocabulary in multilingual models [71, 15, 80, 34] improves performance, especially for low-resource languages. However, large vocabularies [31] increase the computational overhead during both training and generation phases. For example, Liao et al. [35] demonstrate that low-frequency words, due to their limited occurrence in training data, often lack sufficient examples to develop robust representations when vocabularies are excessively large. Dou et al. [19] reveal that expanding vocabularies during continual pre-training can lead to significant performance degradation for low-resource languages. More recently, Dagan et al. [18] explored the trade-offs associated with vocabulary size, proposing optimal vocabulary sizes for both memory efficiency and inference speed in code generation tasks. Our work complements these efforts by focusing on the broader impact of vocabulary size on downstream performance across various tasks. Specifically, we address a critical, under-explored question: How can we optimally allocate vocabulary size to maximize the downstream performance with the same compute budget?